OpenStudy (anonymous):

process of runge kutta method to estimate (0.5) with h=0.25 dy/dx=y-x^2 & y(0)=0.1

OpenStudy (anonymous):

you mean y(0.5)?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

what order?

OpenStudy (anonymous):

4th order sorry my friend had arrived

OpenStudy (anonymous):

required: \[h=.25\quad x_0=0\quad f(x_n)={dy\over dx}=y(x_n)-x_n^2\quad x_N=5\\ N=\frac{x_N-x_0}{h} \] then interationf from n=1 to n=N

OpenStudy (anonymous):

4th order, you have k1, k2, k3, k4 \[k_1=f(x_n,y_n)\\ k_2=f\left(x_n+{h\over2},y_n+{hk_1\over2}\right)\\ k_3=f\left(x_n+{h\over2},y_n+{hk_2\over2}\right)\\ k_4=f(x_n+k,y_n+hk_3) \] and \[ y_{n+1}=y_n+{h\over6}(k_1+2k_2+2k_3+k_4) \]

OpenStudy (anonymous):

start with n=0 to find y_1 for that, you first find the four k and find y_1 then n=1, again find the four k, and find y2 keep going till you get to n=N-1

OpenStudy (anonymous):

oh i see thanks man,you explain this stuff in easy way :)

OpenStudy (ajprincess):

isnt t \(k1=hf(x_n,y_n)\) this?

OpenStudy (anonymous):

oh yeah just checked :) hf()

OpenStudy (anonymous):

I put "h" in evaluating "y" not "k" programatically, this way is efficient for computations

OpenStudy (anonymous):

oh